\begin{thebibliography}{99}
\bibitem{ADG2}
K.\,V. Adaricheva, W. Dziobiak, and V.\,A. Gorbunov,
\emph{Algebraic atomistic lattices of quasivarieties},
Algebra and Logic \textbf{36} (1997), 213--225; translated from
Algebra i Logika \textbf{36} (1997), 363--386.

\bibitem{AGT}
K.\,V. Adaricheva, V.\,A. Gorbunov, and V.\,I. Tumanov,
\emph{Join-semidistributive lattices and convex geometries},
Adv. Math. \textbf{173} (2003), 1--49.

\bibitem{Birk}
G. Birkhoff,
``Lattice Theory'',
7th printing of the 1967 3rd ed., American Mathematical Society Colloquium
Publications \textbf{25}, Providence, Rhode Island: American Mathematical
Society, 1993. vi+418~p.

\bibitem{BB}
G. Birkhoff and M.\,K. Bennett,
\emph{The convexity lattice of a poset}, Order \textbf{2} (1985), 223--242.

\bibitem{CrDi}
P. Crawley and R.\,P. Dilworth,
``Algebraic Theory of Lattices'',
Prentice-Hall, New Jersey, 1973. vi+201~p.

\bibitem{Comp}
G. Gierz, K.\,H. Hofmann, K. Keimel, J.\,D. Lawson, M. Mislove, and D.\,S.
Scott, ``A Compendium of Continuous Lattices'',
Springer-Verlag, Berlin, New York, 1980, xx+371~p.

\bibitem{Gorb}
V.\,A. Gorbunov,
``Algebraic theory of quasivarieties'',
(Algebraicheskaya teoriya kvazimnogoobrazij) (Russian)
Sibirskaya Shkola Algebry i Logiki. 5. Novosibirsk:
Nauchnaya Kniga, 1999. xii+368~p.
English translation by Plenum, New York, 1998. xii+298~p.

\bibitem{GLT2}
G. Gr\"atzer,
``General Lattice Theory. Second edition'', new appendices by the
author with B.\,A. Davey, R. Freese, B. Ganter,
M. Greferath, P. Jipsen, H.\,A. Priestley, H. Rose, E.\,T. Schmidt,
S.\,E. Schmidt, F. Wehrung, and R. Wille. Birkh\"auser Verlag, Basel,
1998. xx+663~p.

\bibitem{Jaku96}
J. Jakubik,
\emph{Directly indecomposable direct factors of a lattice},
Math. Bohem. \textbf{121}, no.~3 (1996), 281--292.

\bibitem{Jano67}
M.\,F. Janowitz,
\emph{The center of a complete relatively complemented lattice is a complete
sublattice}, Proc. Amer. Math. Soc. \textbf{18} (1967), 189--190.

\bibitem{Libk95}
L.~Libkin,
\emph{Direct decompositions of atomistic algebraic lattices},
Algebra Universalis \textbf{33}, no.~1 (1995), 127--135.

\bibitem{Wale97}
A. Walendziak,
\emph{On direct decompositions of lattices},
Algebra Universalis~\textbf{37}, no.~2 (1997), 185--190.

\end{thebibliography}

\documentclass[reqno]{amsart}
\usepackage{amssymb,eucal}

\newcommand{\jz}{\ensuremath{\set{\vee,0}}}
\newcommand{\one}{\mathbf{1}}
\newcommand{\two}{\mathbf{2}}

\newcommand{\jirr}{join-ir\-re\-duc\-i\-ble}

\newcommand{\mirr}{meet-ir\-re\-duc\-i\-ble}
\newcommand{\Mirr}{Meet-ir\-re\-duc\-i\-ble}
\newcommand{\jsd}{join-sem\-i\-dis\-trib\-u\-tive}

\newcommand{\fbs}{finitely bi-spatial}
\newcommand{\Fbs}{Finitely bi-spatial}

\newcommand{\pup}[1]{\textup{(}#1\textup{)}}

\newcommand{\set}[1]{\{#1\}}
\newcommand{\setm}[2]{\{#1\mid#2\}}

\newcommand{\es}{\varnothing}

\newcommand{\id}{\mathrm{id}}

\newcommand{\ol}[1]{\overline{#1}}

\DeclareMathOperator{\Neu}{Neu}
\DeclareMathOperator{\Cen}{Cen}

\newcommand{\SP}{\mathbf{S_p}}
\newcommand{\Co}{\mathbf{Co}}

\DeclareMathOperator{\SPl}{S_p}

\DeclareMathOperator{\J}{J}
\DeclareMathOperator{\M}{M}

\newcommand{\Pow}{\mathcal{P}}
\newcommand{\PO}{\mathcal{P}(\omega)}
\newcommand{\POF}{\mathcal{P}(\omega)/{\mathrm{fin}}}

\theoremstyle{plain}

\newtheorem{lemma}{Lemma}[section]
\newtheorem{theorem}[lemma]{Theorem}
\newtheorem{proposition}[lemma]{Proposition}
\newtheorem{corollary}[lemma]{Corollary}
\newtheorem{example}[lemma]{Example}

\newtheorem*{stat}{\name}
\newcommand{\name}{testing}

\theoremstyle{definition}
\newtheorem{definition}[lemma]{Definition}
\newtheorem*{problem}{Problem}
\newtheorem*{notation}{Notation}

\theoremstyle{remark}
\newtheorem{remark}[lemma]{Remark}

\newcommand{\qedc}{{\qed}~{\rm Claim~{\theclaim}.}}

\begin{document}

\title[Direct decompositions of lattices]
{Direct decompositions of non-algebraic\\
complete lattices}

\author[F.~Wehrung]{Friedrich Wehrung}
\address{CNRS, UMR 6139\\
D\'epartement de Math\'ematiques\\
Universit\'e de Caen\\
14032 Caen Cedex\\
France}
\email{wehrung@math.unicaen.fr}
\urladdr{http://www.math.unicaen.fr/\~{}wehrung}

\subjclass[2000]{06B05, 06B23, 06B35, 06D05}
\keywords{complete, lattice, center, direct product, directly indecomposable,
\jirr, spatial}

\date{\today}

\begin{abstract}
For a given complete lattice $L$, we investigate whether $L$ can be
decomposed as a direct product of directly indecomposable lattices. We prove
that this is the case if every element of $L$ is a join of \jirr\ 
elements and dually, thus extending to non-algebraic lattices a result of
L.~Libkin. We illustrate this by various examples and counterexamples.
\end{abstract}

\maketitle

\section{Introduction}
L. Libkin proves in \cite{Libk95} that if an algebraic lattice $L$ is
\emph{spatial}, that is, every element of $L$ is a join of completely
join-irreducible elements of $L$, then $L$ can be decomposed as a direct
product of directly indecomposable lattices---we say that $L$ is
\emph{totally decomposable}. This result extends the classical one about
decomposing a geometric lattice as a product of indecomposable factors. It
is in turn extended in J.~Jakubik~\cite{Jaku96} by relaxing the
completeness assumptions on $L$, and in A.~Walendziak \cite{Wale97} to
algebraic lattices in which the unit element is a join of \jirr\ elements.
None of these results avoids the assumption that the lattice is compactly
generated, in particular, they do not apply to the closure
lattices of the so-called
\emph{convex geometries} studied in~\cite{AGT}, as the latter are not
algebraic as a rule (by definition, a convex geometry is a
closure space satisfying the anti-exchange property).

In this paper, we extend Libkin's methods and result to a class of
lattices that properly contains both Libkin's lattices and all closure
lattices of most convex geometries, the class of \emph{\fbs} complete
lattices (Definition~\ref{D:finbisp}), see Theorem~\ref{T:DecompL}. 
We also illustrate this by a few examples and counterexamples that show,
in particular, that our assumptions cannot be relaxed much:
\begin{itemize}\em
\item There exists a self-dual, complete, distributive lattice $D$ whose
center is a complete atomistic sublattice but $D$ is not totally
decomposable \pup{see Example~\textup{\ref{Ex:NonDec}}}.

\item There exists a dually algebraic, atomistic, distributive lattice
whose center is not complete \pup{see Example~\textup{\ref{Ex:DCX}}}.

\item Denote by $\SP(A)$ the lattice of algebraic subsets of a complete
lattice $A$. If $A$ is Boolean, then $\SP(A)$ is subdirectly irreducible
\pup{Proposition~\textup{\ref{P:SPIndec}}}, but for $A$ a chain, $\SP(A)$
may not have complete center
\pup{see Example~\textup{\ref{Ex:NonCpleteCen}}}.
\end{itemize}

We observe that Examples~\ref{Ex:DCX} and \ref{Ex:NonCpleteCen} solve
negatively a problem formulated by M.\,F. Janowitz in \cite{Jano67},
whether the center of a complete lattice must be a complete sublattice.

For a set $X$, we denote by $\Pow(X)$ the powerset lattice of $X$. We adopt
the standard set-theoretical notation for ordinals, for example,
$n=\set{0,\ldots,n-1}$ for every nonnegative integer $n$,
then $\omega=\set{0,1,2,\ldots}$, and $\omega+1=\omega\cup\set{\omega}$.

An element $p$ of a lattice $L$ is \emph{\jirr} (resp., \emph{completely
\jirr}), if it is nonzero if $L$ has a zero, and $p=x\vee y$ implies that
$p\in\set{x,y}$, for all $x$, $y\in L$ (resp., $p$ has a unique lower
cover). \Mirr\ (resp., completely \mirr) elements are defined dually.
We denote by $\J(L)$ (resp., $\M(L)$) the set of all \jirr\ (resp., \mirr)
elements of a lattice $L$.

For elements $x$ and $y$ of a given poset, let $x\prec y$ be the statement
that $x<y$ and there is no element strictly between $x$ and $y$.
A lattice $L$ with zero is \emph{atomistic}, if every element of $L$ is a
join of atoms of $L$.

\section{Decompositions of complete lattices}\label{S:DecCpl}

We first recall some standard terminology and facts, see
\cite[Chapter~III, Section~2]{GLT2}.
An element $a$ in a lattice $L$ is \emph{neutral}, if $\set{a,x,y}$
generates a distributive sublattice of $L$, for all $x$, $y\in L$. We shall
denote by $\Neu L$ the subset of all neutral elements of $L$. If $L$ is
\emph{bounded}, we say that an element $a$ of $L$ is \emph{central}, if it is
both neutral and complemented in $L$; then the complement $\neg a$ is
unique, and it is also central. Hence $\Neu L$ is a distributive sublattice
of~$L$, and, if $L$ is bounded, then $\Cen L$ is a Boolean sublattice of
$L$.

The elements of $\Cen L$ correspond exactly to the direct decompositions of
$L$. This can be expressed conveniently in the following way, see
\cite[Theorem~III.4.1]{GLT2}:

\begin{lemma}\label{L:charactC(L)}
Let $L$ be a bounded lattice, let $a$, $b\in L$. Then the following are
equivalent:
\begin{enumerate}
\item There are bounded lattices $A$ and $B$ and an isomorphism
$f\colon L\to A\times B$ such that $f(a)=(1,0)$ and $f(b)=(0,1)$.

\item $(a,b)$ is a complementary pair of elements of $\Cen L$, that is,
$a$, $b\in\Cen L$, $a\nobreak\wedge\nobreak b\nobreak=\nobreak0$, and
$a\vee b=1$.
\end{enumerate}
\end{lemma}

We observe the following easy consequence of Lemma~\ref{L:charactC(L)}:

\begin{proposition}\label{P:At(C(L))}
Let $L$ be a bounded lattice, let $a\in\Cen L$. Then the following assertions
hold:
\begin{enumerate}
\item $\Cen([0,a])=\Cen L\cap[0,a]$.

\item If $a$ is an atom of $\Cen L$, then the interval $[0,a]$ is directly
indecomposable.
\end{enumerate}
\end{proposition}

\begin{definition}
A lattice $L$ is \emph{totally decomposable}, if it is isomorphic to a direct
product of the form $\prod_{i\in I}L_i$, where all the $L_i$-s are directly
indecomposable.
\end{definition}

Totally decomposable complete lattices can be easily characterized as follows:

\begin{proposition}\label{P:TotDec}
Let $L$ be a complete lattice. Then the following are equivalent:
\begin{enumerate}
\item $L$ is totally decomposable;

\item $\Cen L$ is a complete sublattice of $L$, it is atomistic, and, if $U$
denotes the set of its atoms, then the following holds:
 \begin{equation}\label{Eq:Decompx}
 x=\bigvee_{u\in U}(x\wedge u),\quad\text{for all }x\in L.\tag{J}
 \end{equation}
\end{enumerate}
\end{proposition}

\begin{proof}
(i)$\Rightarrow$(ii) Suppose that $L=\prod_{i\in I}L_i$, for a family
$(L_i)_{i\in I}$ of directly indecomposable lattices. Observe that all the
$L_i$-s are complete, in particular, they are bounded lattices. For all
$X\subseteq I$, the characteristic function $\chi_X$ of $X$ in $I$ belongs to
the center of $L$, and its complement is $\chi_{I\setminus X}$. The
complemented pair $(\chi_X,\chi_{I\setminus X})$ of elements of $\Cen L$
induces an isomorphism $L\cong L_X\times L_{I\setminus X}$, where we put
$L_Y=\prod_{i\in Y}L_i$ for every subset $Y$ of $I$. Conversely, if
$u=(u_i)_{i\in I}$ is an element of $\Cen L$, then
$u_i\in\Cen L_i$, for all $i\in I$, thus, since $L_i$ is directly
indecomposable, $u_i\in\set{0,1}$. Therefore, $u=\chi_X$, where
$X=\setm{i\in I}{u_i=1}$.

Consequently, $\Cen L=\setm{\chi_X}{X\subseteq I}$ is a complete sublattice of
$L$. Furthermore, it is atomistic, with atoms the elements $\chi_{\set{i}}$
for $i\in I$. The assertion \eqref{Eq:Decompx} follows easily.

(ii)$\Rightarrow$(i) Suppose that (ii) holds, and denote by $U$ the set of
all atoms of $\Cen L$. Put $L_u=[0,u]$, for all $u\in U$,
then $L'=\prod_{u\in U}L_u$, and define maps
$f\colon L\to L'$ and $g\colon L'\to L$ by the
rules
 \begin{align*}
 f(x)&=(x\wedge u)_{u\in U},&&\text{for all }x\in L,\\
 g\bigl((x_u)_{u\in U}\bigr)&=\bigvee_{u\in U}x_u,&&
 \text{for all }(x_u)_{u\in U}\in L'.
 \end{align*}
For $(x_u)_{u\in U}\in L'$, if we put $x=\bigvee_{u\in U}x_u$,
then, for any $u\in U$, we obtain, by using the fact that $u$ is neutral, the
inequalities $x_u\leq x\wedge u\leq(x_u\vee\neg u)\wedge u=x_u$, whence
$x_u=x\wedge u$. Hence $f\circ g=\id_{L'}$.
Moreover, $g\circ f=\id_L$ follows from the assumption~(ii). Hence, $f$ and
$g$ are mutually inverse isomorphisms. By Proposition~\ref{P:At(C(L))}(ii),
all the factors of the form $L_u$ are directly indecomposable.
\end{proof}

\begin{remark}\label{R:TotDec}
For a bounded lattice $L$, the completeness assumption in
Proposition~\ref{P:TotDec} can be much relaxed. For example,
Proposition~\ref{P:TotDec} remains valid under the assumption that any
family $(x_u)_{u\in U}$ with $x_u\leq u$, for all $u\in U$, has a
join, and the proof is the same.
\end{remark}

In our next result, we shall state a number of conditions that imply
\eqref{Eq:Decompx}. In order to state it conveniently, we set a definition,
that will also be used in Section~\ref{S:Decomp}:

\begin{definition}\label{D:spatial}
Let $L$ be a lattice. We say that $L$ is \emph{finitely spatial} (resp.,
\emph{spatial}), if every element of $L$ is a join of \jirr\ (resp.,
completely \jirr) elements of~$L$. Let \emph{dually spatial}, resp.
\emph{dually finitely spatial}, be the dual notions.
\end{definition}

For example, the real unit interval $[0,1]$ is finitely spatial but not
spatial.

\begin{proposition}\label{P:SuffDec}
Let $L$ be a complete lattice such that $\Cen L$ is a complete atomistic
sublattice of $L$. Then each of the following conditions \pup{and also its
dual} implies that $L$ is totally decomposable:
\begin{enumerate}
\item $L$ is upper continuous.

\item $L$ is \emph{separative}, that is, for any elements $x$, $y\in L$
such that $x\nleq y$, there exists $z\in L$ such that $0<z\leq x$ and
$z\wedge y=0$.

\item $L$ is finitely spatial.
\end{enumerate}
\end{proposition}

\begin{proof}
By Proposition~\ref{P:TotDec}, it suffices that the condition
\eqref{Eq:Decompx} is satisfied by $L$. So let $x\in L$. We put
$y=\bigvee_{u\in U}(x\wedge u)$. In case $L$ is upper continuous, we observe
that $x\wedge\bigvee V=\bigvee_{u\in V}(x\wedge u)$ for every finite subset
$V$ of $U$ (because all elements of $U$ are neutral). Hence, by the upper
continuity of $L$,
 \[
 x=x\wedge\bigvee U=\bigvee_{V\subseteq U\,\text{finite}}
 \left(x\wedge\bigvee V\right)
 =\bigvee_{u\in U}(x\wedge u)=y.
 \]
We conclude the proof of (i) by Proposition~\ref{P:TotDec}.

Suppose that $L$ is separative and that $y<x$. Then, by
assumption, there exists $z\in L$ such that $0<z\leq x$ but $z\wedge y=0$.
Hence, for all $u\in U$, the equality $z\wedge u=0$ holds, thus $z\leq\neg u$.
Therefore, $z\leq\bigwedge_{u\in U}\neg u=\neg\bigvee U=0$, a contradiction.

Finally, suppose that $L$ is finitely spatial. To prove that
\eqref{Eq:Decompx} holds at all elements of $L$, it suffices to verify it for
$x\in\J(L)$. Suppose that it is not the case, that is,
$x>\bigvee_{u\in U}(x\wedge u)$, where $U$ denotes the set of atoms of
$\Cen L$. Every element $u$ of~$U$ belongs to $\Cen L$, whence
$x=(x\wedge u)\vee(x\wedge\neg u)$, but $x\wedge u<x$ by assumption and $x$
is \jirr, thus $x\wedge\neg u=x$, that is, $x\leq\neg u$. This holds for
all $u\in U$, therefore, by assumption on $\Cen L$, $x=0$, a contradiction.
\end{proof}

In particular, we observe that condition (ii) of Proposition~\ref{P:SuffDec}
holds if $L$ is either \emph{atomistic} or \emph{sectionally complemented}.
Since the center of a complete relatively complemented lattice is a complete
sublattice, see \cite{Jano67}, we obtain the following result:

\begin{corollary}\label{C:DecRelCpl}
Let $L$ be a complete relatively complemented lattice. If $\Cen L$ is
atomistic, then $L$ is totally decomposable.
\end{corollary}

To conclude the present section, we shall now see that the condition
\eqref{Eq:Decompx} is not redundant in the statement of
Proposition~\ref{P:TotDec}. 

\begin{example}\label{Ex:NonDec}
There exists a self-dual, complete, distributive lattice $D$ such that
$\Cen D$ is a complete atomistic sublattice of $D$ but $D$ is not totally
decomposable.
\end{example}

\begin{proof}
{}From the classical theory of Boolean algebras, we know that any Boolean
algebra can be embedded into a complete Boolean algebra, see, for example,
\cite[Lemma~II.4.12]{GLT2}. We apply this to the
Boolean algebra $\POF$ of all subsets of $\omega$ modulo the ideal of finite
subsets, to embed it into a complete Boolean algebra $B$. We denote by $[x]$
the equivalence class, modulo the ideal of finite sets, of any subset $x$ of
$\omega$. We observe that $x\mapsto[x]$ defines a homomorphism of Boolean
algebras from $\PO$ to $B$. Thus, the subset $D$ of $\PO\times B\times\PO$
defined as
 \[
 D=\setm{(x,\alpha,y)\in\PO\times B\times\PO}
 {x\subseteq y\text{ and }[x]\leq\alpha\leq[y]}
 \]
is a sublattice of $\PO\times B\times\PO$, in particular, it is a
distributive lattice. Furthermore, $D$ is self-dual, \emph{via} the map
$(x,\alpha,y)\mapsto(\omega\setminus y,\neg\alpha,\omega\setminus x)$.

Let $\varphi\colon\PO\to D$, $x\mapsto(x,[x],x)$. It is obvious that
$\varphi$ is a $0,1$-lattice embedding. Furthermore, since $D$ is a bounded
distributive lattice, the center of $D$ consists exactly of the complemented
elements of $D$. Since $\varphi$ is a $0,1$-lattice homomorphism from $\PO$
to $D$, the range of $\varphi$ is contained in the center of $D$. Conversely,
if $z=(x,\alpha,y)$ is an element of $\Cen D$, then $z$ has a complement,
say, $z'=(x',\alpha',y')\in D$, so $x'=\omega\setminus x$ and
$y'=\omega\setminus y$, thus, since $x\subseteq y$ and $x'\subseteq y'$, we
obtain that $x=y$ and $x'=y'$, whence $\alpha=[x]$, so $z=\varphi(x)$.
Therefore, $\Cen D$ is the range of $\varphi$. It is atomistic, with
atoms the elements $a_n=\varphi(\set{n})=(\set{n},0,\set{n})$, for
$n<\omega$.

We now claim that $D$ is a complete lattice.
Indeed, let $(x_i,\alpha_i,y_i)_{i\in I}$ be a family of elements of~$D$, we
prove that it has a greatest lower bound in $D$. Put $x=\bigcap_{i\in I}x_i$,
$y=\bigcap_{i\in I}y_i$, and $\alpha=\bigwedge_{i\in I}\alpha_i\wedge[y]$. It
is obvious that $(x,\alpha,y)$ belongs to $D$ and that it is contained in
$(x_i,\alpha_i,y_i)$, for all $i\in I$. Let $(x',\alpha',y')\in D$ such
that $(x',\alpha',y')\leq(x_i,\alpha_i,y_i)$, for all $i\in I$. Then
$x'\subseteq x$ and $y'\subseteq y$, thus, since $\alpha'\leq\alpha_i$,
for all $i\in I$, and $\alpha'\leq[y']\leq[y]$, we obtain that
$\alpha'\leq\alpha$. So we have verified that $(x,\alpha,y)$ is the
greatest lower bound of $\setm{(x_i,\alpha_i,y_i)}{i\in I}$ in $D$; whence
$D$ is a complete lattice.

Moreover, in the particular case where $x_i=y_i$, for all $i\in I$ (so
$\alpha_i=[x_i]$), we obtain that $(x,\alpha,y)=(x,[x],x)$, where
$x=\bigcap_{i\in I}x_i$. Hence, $\varphi$ is a complete meet embedding.
The verification of the fact that $\varphi$ is a complete join embedding is
similar. Hence, $\varphi$ is a complete lattice embedding from $\PO$ into
$D$. Therefore, the center of $D$, which is also the range of $\varphi$, is a
complete sublattice of $D$.

Now put $b=(\es,1,\omega)$ (so $b\in D$).
We observe that $b\wedge a_n=(\es,0,\set{n})$, for all $n<\omega$,
hence
 \[
 \bigvee_{n<\omega}(b\wedge a_n)=(\es,0,\omega)<b.
 \]
By Proposition~\ref{P:TotDec}, $D$ is not totally decomposable.
\end{proof}

\begin{remark}
It is easy to read, in the proof above, the places where
Example~\ref{Ex:NonDec} fails the conditions (i)--(iii) of
Proposition~\ref{P:SuffDec}. For all $n<\omega$, the element
$\ol{a}_n=(n,0,n)$ belongs to $D$, while $\bigvee_{n<\omega}\ol{a}_n=1$ and
$\bigvee_{n<\omega}(\ol{a}_n\wedge b)<b$, thus verifying that $D$ is not
upper continuous. Put $\ol{b}=(\es,0,\omega)$. Then $\ol{b}<b$, while there is no
nonzero $z\leq b$ such that $z\wedge\ol{b}=0$, thus verifying that $D$ is
not separative. Finally, the \jirr\ elements below $b$ are exactly all the
$(\es,0,\set{n})$, and these join to $\ol{b}<b$, thus verifying that $D$
is not finitely spatial.

\end{remark}

\section{\Fbs\ complete lattices}
\label{S:Decomp}

We start by defining the objects of the section title:

\begin{definition}\label{D:finbisp}
We say that a bounded lattice $L$ is \emph{\fbs}, if it is both finitely
spatial and dually finitely spatial (see Definition~\ref{D:spatial}).
\end{definition}

\begin{notation}
Let $x\in L$, let $(x_i)_{i\in I}$ be a family of elements
of $L$. Let $x=\bigvee^*_{i\in I}x_i$ hold, if
 \[
 p\leq x\quad\text{if{f}}\quad\exists i\in I\text{ such that }
 p\leq x_i,\text{ for all }p\in\J(L).
 \]
For $|I|=2$, we define similarly the notation $z=x\vee^*y$, for $x$, $y$,
$z\in L$. Similarly, let $x=\bigwedge^*_{i\in I}x_i$ hold, if
 \[
 x\leq u\quad\text{if{f}}\quad\exists i\in I\text{ such that }
 x_i\leq u,\text{ for all }u\in\M(L),
 \]
and, for $|I|=2$, we define similarly the notation $z=x\wedge^*y$.
\end{notation}

The following lemma is similar in essence to \cite[Lemma~1]{Libk95}:

\begin{lemma}\label{L:N(L)commutes}
Let $L$ be a bounded lattice, let $a\in\Neu L$, let $x$, $y\in L$. Then
$y=a\vee x$ \pup{resp., $y=a\wedge x$} implies that $y=a\vee^*x$
\pup{resp., $y=a\wedge^*x$}.
\end{lemma}

\begin{proof}
We prove, for example, that $y=a\vee x$ implies that $y=a\vee^*x$. Let
$p\in\J(L)$ such that $p\leq y$. Then, by using the fact that $a$ is neutral,
$p=p\wedge(a\vee x)=(p\wedge a)\vee(p\wedge x)$, hence, since $p$ is \jirr,
either $p\leq a$ or $p\leq x$. The proof for the meet is similar.
\end{proof}

We leave to the reader the straightforward proof of the following lemma:

\begin{lemma}\label{L:sqtononsq}
Let $L$ be a \fbs\ bounded lattice.
Let $x$, $y\in L$, let $(x_i)_{i\in I}$ be a family of elements of $L$. Then
the following assertions hold:
\begin{enumerate}
\item $x=\bigvee^*_{i\in I}x_i$ implies that $x=\bigvee_{i\in I}x_i$;

\item $x=\bigwedge^*_{i\in I}x_i$ implies that $x=\bigwedge_{i\in I}x_i$;

\item $x=\bigvee^*_{i\in I}x_i$ implies that
$x\wedge y=\bigvee^*_{i\in I}(x_i\wedge y)$;

\item $x=\bigwedge^*_{i\in I}x_i$ implies that
$x\vee y=\bigwedge^*_{i\in I}(x_i\vee y)$.
\end{enumerate}
\end{lemma}

\begin{lemma}\label{L:CplPair}
Let $L$ be a \fbs\ bounded lattice.
Let $a$, $b\in L$. Then the following are equivalent:
\begin{enumerate}
\item $a\vee^*b=1$ and $a\wedge^*b=0$;

\item $(a,b)$ is a complementary pair of elements of $\Cen L$.
\end{enumerate}

\end{lemma}

\begin{proof}
(i)$\Rightarrow$(ii) We consider the maps $f\colon L\to[0,a]\times[0,b]$ and
$g\colon[0,a]\times[0,b]\to L$ defined by the following formulas:
 \begin{align*}
 f(z)&=(z\wedge a,z\wedge b),&&\text{for all }z\in L,\\
 g(x,y)&=x\vee y,&&\text{for all }(x,y)\in[0,a]\times[0,b].
 \end{align*}
For any $z\in L$, it follows from Lemma~\ref{L:sqtononsq} that
$z=(z\wedge a)\vee(z\wedge b)$, so
$g\circ f=\id_L$. Conversely, let $x\leq a$ and $y\leq b$ in $L$. Then, again
by using Lemma~\ref{L:sqtononsq},
$x\leq(x\vee y)\wedge a\leq(x\vee b)\wedge(x\vee a)=x\vee(a\wedge^*b)=x$,
whence $x=(x\vee y)\wedge a$. Similarly, $y=(x\vee y)\wedge b$. Therefore,
$f\circ g=\id_{[0,a]\times[0,b]}$, so $f$ and $g$ are mutually
inverse isomorphisms. The conclusion (ii) follows then from
Lemma~\ref{L:charactC(L)}.

(ii)$\Rightarrow$(i) follows immediately from Lemma~\ref{L:N(L)commutes}.
\end{proof}

Now we can prove one of the main lemmas of this section:

\begin{lemma}\label{L:CplfbsC(L)}
Let $L$ be a \fbs\ complete lattice. Then the center $\Cen L$ is a complete
sublattice of $L$.
\end{lemma}

\begin{proof}
Let $(a_i)_{i\in I}$ be a family of elements of $\Cen L$, then
put $a=\bigvee_{i\in I}a_i$ and $b=\bigwedge_{i\in I}\neg a_i$.

We first claim that $a\vee^*b=1$ and $a\wedge^*b=0$. Indeed, let us prove
for example the first assertion. Let $p\in\J(L)$. It follows from 
Lemma~\ref{L:CplPair} that for all $i\in I$, either $p\leq a_i$ or
$p\leq\neg a_i$. Hence, if $p\nleq b$, then there exists $i\in I$ such that
$p\nleq\neg a_i$, whence $p\leq a_i\leq a$. The proof of $a\wedge^*b=0$ is
dual. It follows, again by Lemma~\ref{L:CplPair}, that $(a,b)$ is a
complementary pair of $\Cen L$. In particular, $\Cen L$ is a complete
sublattice of $L$.
\end{proof}

By Lemma~\ref{L:CplfbsC(L)}, for any $x\in L$, there is a least element
$u$ of $\Cen L$ such that $x\leq u$, we denote this element by $e(x)$, the
\emph{central cover} of $x$.

\begin{lemma}\label{L:C(L)atomic}
Let $L$ be a \fbs\ complete lattice.
The Boolean lattice $\Cen L$ is atomistic, with atoms the $e(p)$ for
$p\nobreak\in\nobreak\J(L)$.
\end{lemma}

\begin{proof}
Observe first the obvious equality
$u=\bigvee\setm{e(p)}{p\in\J(L),\ p\leq u}$,
for any $u\in\Cen L$. Hence, it suffices to prove that $e(p)$ is
an atom of $\Cen L$, for all $p\in\J(L)$. Suppose otherwise. Then
$e(p)=u\vee v$, for nonzero elements $u$ and $v$ of $\Cen L$ such that
$u\wedge v=0$. {}From $u\in\Neu L$ follows that
$p=p\wedge(u\vee v)=(p\wedge u)\vee(p\wedge v)$, whence, since $p\in\J(L)$,
either $p\leq u$ or $p\leq v$. Suppose, for example, that $p\leq u$. Then
$e(p)\leq u$, whence $v=0$, a contradiction.
\end{proof}

{}From Proposition~\ref{P:SuffDec}(iii) and Lemmas \ref{L:CplfbsC(L)} and
\ref{L:C(L)atomic}, we can now deduce immediately the main result of this
section:

\begin{theorem}\label{T:DecompL}
Every \fbs\ complete lattice is isomorphic to a direct product of directly
indecomposable lattices.
\end{theorem}

As immediate corollaries of Theorem~\ref{T:DecompL} and the fact that every
algebraic lattice is dually spatial (see \cite[Theorem I.4.22]{Comp}, or
\cite[Lemma~1.3.2]{Gorb}), we observe the following, see
\cite[Theorem~2]{Libk95}:

\begin{corollary}[Libkin's Decomposition Theorem]
\label{C:Libkin}
Every algebraic and spatial lattice is isomorphic to a direct product of
directly indecomposable lattices.
\end{corollary}

In particular, every algebraic and atomistic lattice is isomorphic to a
direct product of directly indecomposable lattices. In fact, since every
algebraic lattice is dually spatial, Theorem~\ref{T:DecompL} makes it
possible to extend Corollary~\ref{C:Libkin} to finitely spatial algebraic
lattices. In particular, we obtain the following consequence, a stronger
form of which is stated in \cite[Corollary~2]{Wale97}:

\begin{corollary}\label{C:AtduAt}
Every algebraic and dually algebraic lattice is isomorphic to a direct
product of directly indecomposable lattices.
\end{corollary}

\begin{example}\label{Ex:DCX}
There exists a dually algebraic, atomistic, distributive lattice $D$ whose
center $\Cen D$ is not complete. In particular, $D$ cannot be decomposed as
a direct product of directly indecomposable lattices.
\end{example}

\begin{proof}
We recall that the \emph{interval topology} on a \emph{totally} ordered set
$T$ is the least topology on $T$ for which all intervals of the form $[a)$
(resp., $(a]$) are closed subsets. It is a well-known result, due to
O.~Frink (see for example \cite[Theorem~X.12.20]{Birk}), that states
that the interval topology on $T$ is compact Hausdorff if{f} $T$ is a
complete lattice.

Now we endow the ordinal $\omega+1$ with its interval topology, and we let
$D$ be the lattice of all closed sets of this topology. Hence,
 \[
 D=\setm{x\subset\omega}{x\text{ is finite}}\cup
 \setm{x\cup\set{\omega}}{x\subseteq\omega}.
 \]
Observe that $D$ is a closure system in the powerset algebra $\Pow(\omega+1)$
of $\omega+1$, thus it is a complete lattice. Moreover, $D$ is a distributive
sublattice of $\Pow(\omega+1)$, and it is atomistic since every element of $D$
is a union of singletons. Moreover, it is straightforward to compute that
 \[
 \Cen D=\setm{x\subset\omega}{x\text{ is finite}}\cup
 \setm{x\cup\set{\omega}}{x\subseteq\omega\text{ is cofinite}},
 \]
so $\Cen D$ consists exactly of the \emph{clopen} subsets of $\omega+1$.
Since $\omega+1$ is a compact topological space, every element of $\Cen D$
is dually compact in $D$. Furthermore, every closed subset of $\omega+1$
is an intersection of clopen subsets, therefore, \emph{$D$ is dually
algebraic}.

Put $a=\setm{2m+1}{m<\omega}$ and $b_n=(\omega+1)\setminus\set{2n}$, for
all $n<\omega$. Observe that both $a\cap m$ and $b_n$ belong to $\Cen D$,
for all $m$, $n<\omega$, and that $a\cap m\subset b_n$. However, there is
no element $x$ of $\Cen D$ such that $a\cap m\subseteq x\subseteq b_n$ for
all $m$, $n<\omega$, because otherwise either $x=a$ or
$x=a\cup\set{\omega}$ would belong to $\Cen D$, a contradiction.
\end{proof}

\begin{remark}
It is easy to verify that $D$ is even \emph{strongly atomic}, that is,
$a<b$ implies that there exists $x\in D$ such that $a\prec x\leq b$, for all
$a$, $b\in D$. We recall that every algebraic
lattice $A$ is \emph{weakly atomic}, that is, for all $a<b$ in $A$, there
are $x$, $y\in A$ such that $a\leq x\prec y\leq b$ (see
\cite[Lemma~2.2]{CrDi} or \cite[Exercise~1.3.1]{Gorb}).
\end{remark}

\section{Direct decompositions of lattices of algebraic subsets}
\label{S:AlgSubs}

For a complete lattice $A$, a subset $X$ of $A$ is \emph{algebraic}, if $X$
is closed under arbitrary intersections and nonempty up-directed joins, and we
denote by $\SP(A)$ the lattice of all algebraic subsets of $A$. Then the
following basic lemma holds, see~\cite{AGT} for more information:

\begin{lemma}
Let $A$ be a complete lattice. Then the following assertions hold:
\begin{enumerate}
\item If $A$ is upper continuous, then $\SP(A)$ is a \jsd, lower continuous
lattice.

\item If $A$ is algebraic, then $\SP(A)$ is dually algebraic.

\item If $A=\Pow(X)$ for some set $X$, then $\SP(A)$ is dually spatial.
\end{enumerate}
\end{lemma}

We recall at this point that any algebraic lattice is upper continuous (see
\cite[Lemma~2.3]{CrDi}), and that for $A$ a general algebraic lattice,
$\SP(A)$ does not need to be dually spatial (see \cite{AGT}). We also observe
that $\SP(A)$ is the closure lattice of the atomistic closure space
$(A\setminus\set{1},\SPl)$, where, for every subset $X$ of $A$,
we put $\SPl(X)=\ol{X}\setminus\set{1}$, where $\ol{X}$ denotes the
algebraic subset of $A$ generated by $X$.

K.\,V. Adaricheva has kindly informed the author that all lattices of the
form $\SP(\Pow(X))$ are directly indecomposable. A stronger result is the
following:

\begin{proposition}\label{P:SPIndec}
For any complete Boolean algebra $B$, the lattice $\SP(B)$ is
subdirectly irreducible.
\end{proposition}

\begin{proof}
The atoms of $\SP(B)$ are the $U_a=\set{a,1}$ for
$a\in B\setminus\set{1}$. For $X$, $Y\in\SP(B)$,
we denote by $\Theta(X,Y)$ the principal congruence of
$\SP(B)$ generated by the pair $(X,Y)$.
For any $a\in B\setminus\set{0,1}$, the containment
$U_0\subset U_a\vee U_{\neg a}$ holds, with $U_0$, $U_a$, and $U_{\neg a}$
distinct atoms of $\SP(B)$, thus
$\Theta(\set{1},U_0)\subseteq\Theta(\set{1},U_a)$.
Since $\SP(B)$ is atomistic, it follows that $\Theta(\set{1},U_0)$ is the
smallest nonzero congruence of~$\SP(B)$.
\end{proof}

\begin{remark}\label{Rk:Notsimple}
The lattice $\SP(\Pow(2))$ is the (finite) \jz-semilattice defined by
generators $a$, $b$, $c$ and the unique relation $c\leq a\vee b$, hence it is
not simple.
\end{remark}

\begin{remark}\label{Rk:NotSI}
Even for finite atomistic lattices which are lower bounded homomorphic
images of free lattices, direct indecomposability is not equivalent to
subdirect irreducibility. For example, the lattice $\Co(\two^2)$ of all
convex subsets of $\two^2$ (diagrammed for example in \cite[p.~224]{BB})
is directly indecomposable, although not subdirectly irreducible. This is
another strong point of contrast between geometric lattices and convex
geometries (see \cite{AGT} for the latter): namely, every directly
indecomposable geometric lattice is subdirectly irreducible, see
\cite[Theorem~IV.3.6]{GLT2}.
\end{remark}

On the other hand, as we shall see in a moment, the lattice $\SP(A)$ displays
a very different behavior for $A$ a \emph{totally ordered} algebraic lattice.
The proof of the following lemma is a straightforward exercise.

\begin{lemma}\label{L:TopAlgClos}
Let $A$ be a totally ordered algebraic lattice. Then a subset $X$ of $A$
belongs to $\SP(A)$ if{f} $X$ is closed for the interval topology and
$1\in X$.
\end{lemma}

\begin{example}[see \cite{ADG2}]\label{Ex:NonCpleteCen}
Let $C=[0,1]$ be the rational unit interval, let $A$ be the ideal lattice of
$C$. Then $\Cen\SP(A)$ is not a complete lattice.
\end{example}

\begin{proof}
Put $j(x)=[0,x]$, for all $x\in C$, and, if $x>0$, put $j(x)_*=[0,x)$,
so $j(x)_*\prec j(x)$.
Observe that $A=\setm{j(x)}{x\in C}\cup\setm{j(x)_*}{x\in C\setminus\set{0}}$
is a complete chain with top element $\one=[0,1]$. It follows from
Lemma~\ref{L:TopAlgClos} that $\SP(A)$ is isomorphic to the lattice $D$ of
all closed subsets of $A\setminus\set{\one}$ endowed with the interval
topology. Therefore, the center $\Cen\SP(A)$ is isomorphic to the Boolean
lattice $B$ of all clopen subsets of $A\setminus\set{\one}$ for the interval
topology.

Now put $a_n=\frac{1}{2}-\frac{1}{2n}$ for every positive integer $n$, and
$a=\frac{1}{2}$. For each positive integer $n$, we put
 \[
 X_n=[j(a_{2n}),j(a_{2n+1})_*],\qquad
 Y_n=\bigcup_{0<k\leq n}X_k\cup[j(a_{2n+2}),j(a)_*].
 \]
Then both $X_m$ and $Y_n$ are clopen subsets of
$A\setminus\set{\one}$ with $X_m\subset Y_n$, for all $m$, $n>0$. However,
the only subsets $Y$ of $A\setminus\set{\one}$ such that
$X_m\subseteq Y\subseteq Y_n$ for all $m$, $n>0$ are
$Z=\bigcup_{0<k<\omega}X_k$, which is not closed, and $Z\cup\set{j(a)_*}$,
which is not open.
\end{proof}

We conclude the paper with a problem:

\begin{problem}
Find a common generalization of Theorem~\ref{T:DecompL} (decomposition
theorem for finitely bi-spatial complete lattices) and various
decomposition results such as the ones in \cite{Jaku96,Jano67,Wale97}.
\end{problem}

Indeed, the hard core of Theorem~\ref{T:DecompL} and its analogues lies in
proving that the center is complete. All the methods used here and in
\cite{Jaku96,Jano67,Wale97} bear some formal similarity, but none of the
results seems to follow from the others.

\section*{Acknowledgment}
The author is grateful to Kira Adaricheva for having gotten him
interested in the topic and for her many helpful comments.

\end{document}